# Computing the number of realisations of a rigid graph

### v1

@@ Appendix A The degree of a graph map is the realisation number of a graph
Since $G$ is rigid, $p_{G}$ is generically finite. Choose a generic point $\lambda\in I_{G}$. By Lemma 14(i), this is equivalent to choosing a generic realisation $\rho\in\mathbb{C}^{2|V(G)|}$ and fixing $\lambda=f_{G}(\rho)$. Fix $p_{G}^{-1}(\lambda)=\{\mu_{1},\ldots,\mu_{s}\}$ and fix $C_{1},\ldots,C_{t}$ to be the connected components of $f^{-1}_{G}(\lambda)$. By Lemma 14(i) and [JO19, Theorem 3.6], $t=2c_{2}(G)$, with each connected component consisting of the orbit of a single realisation under the orientation-preserving [^2] Every affine map $E_{2}$ is of the form $z\mapsto T(z)+z_{0}$ for some vector $z_{0}\in\mathbb{C}^{2}$ and some linear $2\times 2$ matrix $T$ where $T^{T}T=TT^{T}=I_{2}$. Either $\det(T)=1$ and $T$ is orientation-preserving, or $\det(T)=-1$ and $T$ is not orientation-preserving. affine transformations of $E_{2}$; moreover, if we set $c=c_{2}(G)$, we can relabel our connected components $C_{\pm 1},\ldots,C_{\pm c}$ so that $C_{-i}$ is the image of $C_{i}$ under the reflection $(x,y)\mapsto(x,-y)$.
For each $i\in\{\pm 1,\ldots,\pm c\}$, set $\sigma(i)\in\{1,\ldots,s\}$ to be the index such that $f_{K_{V(G)}}(C_{i})=\{\mu_{\sigma(i)}\}$. It is immediate that $\sigma(-i)=\sigma(i)$. As each map in the above diagram is dominant and $p_{G}$ is generically finite, the genericity of $\lambda$ implies that the map $\sigma$ is surjective (and so $s\leq 2c$). Suppose that there exists $i,j\in\{1,\ldots,c\}$ such that $\sigma(i)=\sigma(j)$. This implies the existence of realisations $\rho_{i},\rho_{j}\in f_{G}^{-1}(\lambda)$ where $f_{K_{V(G)}}(\rho_{i})=f_{K_{V(G)}}(\rho_{j})$ but such that $h(\rho_{i})\neq h(\rho_{j})$ for each $h\in E_{2}$. However, this now directly contradicts Lemma 14(ii). Hence, $s=2c$, which implies the desired result. ∎
@@ Appendix B Graph encodings
In the following (Tables 4 and 5) we provide graphs which obtain the realisations numbers that are used in computational results (Section 5).
@@
Table 4: Certificate graphs for the realisation numbers in Table 2.
@@
Table 5: Certificate graphs for the realisation numbers in Table 3.

### v2

@@ Appendix A The degree of a graph map is the realisation number of a graph
Since $G$ is rigid, $p_{G}$ is generically finite. Choose a generic point $\lambda\in I_{G}$. By Lemma˜14(i), this is equivalent to choosing a generic realisation $\rho\in\mathbb{C}^{2|V(G)|}$ and fixing $\lambda=f_{G}(\rho)$. Fix $p_{G}^{-1}(\lambda)=\{\mu_{1},\ldots,\mu_{s}\}$ and fix $C_{1},\ldots,C_{t}$ to be the connected components of $f^{-1}_{G}(\lambda)$. By Lemma˜14(i) and [JO19, Theorem 3.6], $t=2c_{2}(G)$, with each connected component consisting of the orbit of a single realisation under the orientation-preserving [^2] Every affine map $E_{2}$ is of the form $z\mapsto T(z)+z_{0}$ for some vector $z_{0}\in\mathbb{C}^{2}$ and some linear $2\times 2$ matrix $T$ where $T^{T}T=TT^{T}=I_{2}$. Either $\det(T)=1$ and $T$ is orientation-preserving, or $\det(T)=-1$ and $T$ is not orientation-preserving. affine transformations of $E_{2}$; moreover, if we set $c=c_{2}(G)$, we can relabel our connected components $C_{\pm 1},\ldots,C_{\pm c}$ so that $C_{-i}$ is the image of $C_{i}$ under the reflection $(x,y)\mapsto(x,-y)$.
For each $i\in\{\pm 1,\ldots,\pm c\}$, set $\sigma(i)\in\{1,\ldots,s\}$ to be the index such that $f_{K_{V(G)}}(C_{i})=\{\mu_{\sigma(i)}\}$. It is immediate that $\sigma(-i)=\sigma(i)$. As each map in the above diagram is dominant and $p_{G}$ is generically finite, the genericity of $\lambda$ implies that the map $\sigma$ is surjective (and so $s\leq 2c$). Suppose that there exists $i,j\in\{1,\ldots,c\}$ such that $\sigma(i)=\sigma(j)$. This implies the existence of realisations $\rho_{i},\rho_{j}\in f_{G}^{-1}(\lambda)$ where $f_{K_{V(G)}}(\rho_{i})=f_{K_{V(G)}}(\rho_{j})$ but such that $h(\rho_{i})\neq h(\rho_{j})$ for each $h\in E_{2}$. However, this now directly contradicts Lemma˜14(ii). Hence, $s=2c$, which implies the desired result. ∎
@@ Appendix B Graph encodings
In the following (Tables˜4 and 5) we provide graphs which obtain the realisations numbers that are used in computational results (Section˜5).
@@
Table 4: Certificate graphs for the realisation numbers in Table˜2.
@@
Table 5: Certificate graphs for the realisation numbers in Table˜3.